Aug 20, 2020 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Time: 11:00 a.m. JLab time

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

  1. Short term
    1. Richard Jones' work on G3vG4: run photon, pi+, pi-, proton particle guns. Who?
    2. Efficiency study: model mass and t-distributions
    3. Hadronic split off studies
    4. Covariance matrices for photons; should be redone.
    5. BCAL Wiki reorganization -> All
  2. Long term items
  3. Complete Calorimeter work packages

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run Update
    1. FCAL
    2. BCAL
  4. Monitoring
    1. BCAL LED
      1. BCAL LED Monitoring fall 2018 (Sarah) [2]
      2. Dark rate (SiPM) study update (Varun) [3]
    2. BCAL (offline)
      1. Brief monitoring update, batches 5-8 "redo" (Tolga) [4]
    3. FCAL (offline) FCAL monitoring V (Su)
  5. Calibration
  6. Efficiency
  7. Reconstruction
  8. Simulations
    1. Log Entry 3830769 Hadronic Efficiencies (Elton)
      • CCDB: BCAL/channel_mc_efficiency
      • M22U S3 L4, eff=0 R72120-72435
      • M37D S4 L1, eff=0 R72316-72369
  9. Any other business


Attending: Varun, Sarah, Zisis, Tolga, Dhruval, Cplin, Mark, Karthik, Matt, Rebecca, Elton, Jon, Sean, Susan

  1. Announcements: none
  2. Action Items
    1. Short Term
      1. Particle guns: Regina will do.
      2. Efficiency: See agenda below
      3. Splitoffs: Matt reported that Rebecca will check the neural network on data and MC to see if it is modeled correctly (this was done based on data). This is connected to the Task Force on photon reconstruction. It is a short term job. Zisis mentioned that a new MSc student (Azizah Mahmood) will likely work on ML for FCAL-II. Matt mentioned that any work with MC sims needs to be validated with data.
      4. Covariance matrices: nothing to report
      5. Wiki: Colin will start the FCAL Wiki reorganization.
    2. Long Term: nothing to report.
    3. Work packages: not discussed today.
  3. Run Update
    1. FCAL:
      1. Bases: Mark reported that all spare bases have been used in replacing faulty ones placing the FCAL in the green. Another 20 or so will be needed in the next couple of weeks. Matt suggested that the PMTs+Bases be sorted and classified based on the particular problem (lemo cable connections, coax ripped out of socker, high current draw). Some of these can be repaired by Chris at the lab, others require the application of conformal coating so Paul Smith should be consulted. Colin added that plans should be made during the shutdown to be proactive and accommodate PrimEx requirements with good PMT+Base in the inner rings.
      2. Scalers: two channels w/o scaler rate but occupancy looks normal. Check on the scope; Mark will do at next access.
    2. BCAL: nominal operation. Elton asked about the FP firing. Varun is changing the code to use energy as a criterion (used to be like this when Foda first designed it) instead of the number of hits being > 1200 to indicate an LED event. [After the meeting Varun checked in the code and Mark pulled it.] This will help us identify LED pulses that are stolen by the physics trigger.
  4. Monitoring
    1. BCAL LED
      1. BCAL LED Monitoring fall 2018 (Sarah Shoobert): Motivated by features seen in LEDs and SIPMs in the fall 2019 data (e.g slope versus in run number), Sarah carried out a study on the fall 2018 LED events. Two SiPMs were identified as having an issue (1118 was known, 1405 the new one) whereas many LEDs are showing fluctuations. The full list is in her slides. Fall 2018 overall exhibits same bahaviour as Spring 2019.
      2. Dark rate (SiPM) study update (Varun) [5]: An update was given by Varun (working with Jon), who plotted the pedestal RMS. On slide 3 we can see that above 1 V overbias, the SiPM dominates but below the flash is causing fluctuations. Elton pointed out that their size is small: 1/100 of an ADC count. Mark added that with the flash limitations we can only set an upper bound in extracting the SiPM radiation damage. Varun also showed that channel 1486 shows a drop of 7% in gain in the current (July-August 2020) run. It is the only one that shows such behaviour.
    2. BCAL (offline)
      1. Monitoring update, batches 5-8 "redo" (Tolga): Things are largely ok with some minor issues, most known. Recon2 is better in v16 over v12. Karthik asked about the situation with channel 1118 (see slide 3). Elton replied that it is deadened in the MC but not yet the data. It was agree it to zero it in the ccdb gains table for the runs affected. Later, we may consider adding a dedicated table for dead channels like was done for other detectors. Sean expessed concern over the timing issues on slide 5; Mark will examine the low-lelvel timing to investigate further.
    3. FCAL (offline) (Susan): Susan was on remote shift so asked the group to look at her slides offline and send her comments.
  5. Calibration: Karthik reported that he is running new jobs with different pixel/count factors to re-examine the nonlinearity correction, because the last calibration brought out the pi0 mass 2% higher than expected.
  6. Efficiency costheta (Elton): An update on the costheta efficiency was given by Elton, in connection to the Task Force work. This is a relative method and assumes that inefficiencies are due to the low energy photon. -t slope > 0.2 cut was used for the data. Fall 2018 run period is closer to fall 2017. Slide 11 is the money plot: data vs MC: BCAL is flat for Eg>0.2 GeV. FCAL looks good too. FCAL-BCAL has a bump at 0.2 GeV (because of BCAL behaviour). Note that the method is relative and complex, and therefore hard to drill down to see what is actually happening. MC is complex especially in comparing G3 (HADR flag) to G4, which can affect things as the analysis includes pi+. pi- and the proton. Sean asked about the covariance matrices; once these are available the analysis will be looked at again.
  7. Reconstruction: nothing to report.
  8. Simulations
    1. Hadronic Efficiencies (Elton): Elton looked at suspicious/problematic channels, comparing to observations by Tolga and Varun. A question was asked about the index (which labels the channel); e.g. index 97 corresponds to channel 1118. Karthik will check the indices used by Elton, Varun and himself to ensure they are consistent. Channel 1118 has been deadened in MC and soon will be so in data too, for affected runs.
  9. Any other business: none.